how would you...
A.) use the distributive property to solve and create a binomial to set up the FOIL method
B.) Solve with the FOIL method
could you break it down step by step?
4x(x^2-2x+5)
: how would you...
A.) use the distributive property to solve and create a binomial to set up the FOIL method
B.) Solve with the FOIL method
could you break it down step by step?
4x(x^2-2x+5)

Answer by stanbon(19020)

A.) use the distributive property to solve and create a binomial to set up the FOIL method
(x+a)(x+b)
= (x+a)x + (x+a)b
= xx + ax + bx + ab
Which is F--O--I--L
================================
B.) Solve with the FOIL method
could you break it down step by step?
= 4x(x^2-2x+5)
= 4x*x^2 -4x*2x + 4x*5
============================
Cheers,
Stan H.
Question 151323: how would you...
A.) use the distributive property to solve and create a binomial to set up the FOIL method
B.) Solve with the FOIL method
could you break it down step by step?
4x(x^2-2x+5)
: how would you...
A.) use the distributive property to solve and create a binomial to set up the FOIL method
B.) Solve with the FOIL method
could you break it down step by step?
4x(x^2-2x+5)

Answer by jim_thompson5910(9402)

This problem does not involve FOILing, but involves distribution (you use FOIL when you have two binomials). So I would double check the problem.


Anyways, here's how you would simplify the problem.


4x(x^2-2x+5) Start with the given expression.


4x(x^2)+4x(-2x)+4x(5) Distribute.


4x^3+4x(-2x)+4x(5) Multiply 4x and x^2 to get 4x^3


4x^3-8x^2+4x(5) Multiply 4x and -2x to get -8x^2


4x^3-8x^2+20x Multiply 4x and 5 to get 20x


So 4x(x^2-2x+5) distributes and simplifies to 4x^3-8x^2+20x


In other words, 4x(x^2-2x+5)=4x^3-8x^2+20x